OpenStudy (anonymous):

Solve the system:y-3x=5 y+x=3 write your answer as an ordered pair (x,y). Change fractions to decimals if needed.

OpenStudy (saifoo.khan):

So where are you stuck?

OpenStudy (anonymous):

i dont understand how to do it at all..... i was having surgery when they were teaching it......

OpenStudy (anonymous):

i have about 9 more like it to... so i need to understand how to do it...

OpenStudy (anonymous):

so i know you have to make one of the y's a negative so you can cancel them out and find out x

OpenStudy (anonymous):

so i multiplied the first one by -1 and got -y+3x=-5 y+x=3 so the y's cancel each other out now than you add everything so it is 4x=-1 and right here i get stuck i dont know how to get the x alone

OpenStudy (saifoo.khan):

Sorry i was in another tab.

OpenStudy (saifoo.khan):

You just did 1 mistake. -5 + 3 = -2 You wrote -1

OpenStudy (anonymous):

oh oops...

OpenStudy (anonymous):

so how do you get the x alone now

OpenStudy (saifoo.khan):

so it is 4x=-2 x = -2/4 x = -1/2 Agree?

OpenStudy (anonymous):

wait so you just divide the 4 with the -2 to get the x alone? correct

OpenStudy (saifoo.khan):

Yes sir.

OpenStudy (anonymous):

wait i got 1 then..

OpenStudy (saifoo.khan):

1? Where?

OpenStudy (anonymous):

wait isn't 4/-2 =1

OpenStudy (saifoo.khan):

it's -2/4 not 4/-2 And btw 4/-2 is -2.

OpenStudy (anonymous):

haha ok thank you tons!

OpenStudy (saifoo.khan):

Lol. You're welcome. :)